# LangGPT: Rethinking Structured Reusable Prompt Design Framework for LLMs from the Programming Language

### v1

@@ Abstract
LLMs have demonstrated commendable performance across diverse domains. Nevertheless, formulating high-quality prompts to effectively instruct LLMs poses a challenge for non-AI experts. Existing research in prompt engineering suggests somewhat fragmented optimization principles and designs empirically dependent prompt optimizers. Unfortunately, these endeavors lack a structured design template, incurring high learning costs and resulting in low reusability. Inspired by structured reusable programming languages, we propose LangGPT, a dual-layer prompt design framework as the programming language for LLMs. LangGPT has an easy-to-learn normative structure and provides an extended structure for migration and reuse. Experiments illustrate that LangGPT significantly enhances the capacity of LLMs to produce responses of superior quality compared to baselines. Moreover, LangGPT has proven effective in guiding LLMs to generate high-quality prompts. We have built a community on LangGPT to facilitate the tuition and sharing of prompt design. We also analyzed the ease of use and reusability of LangGPT through a community user survey.
@@ 1 Introduction
Large language models (LLMs) like ChatGPT Achiam et al. (2023); Schulman et al. (2023) can execute diverse tasks Sun et al. (2023b, c); Yu et al. (2023) based on powerful language comprehension, reasoning, and generation capabilities. Injecting domain knowledge also enables LLMs to perform domain-related specific tasks Wang et al. (2023a); Li et al. (2023b); Zhang et al. (2023); Ren et al. (2023). Fully unleashing these capabilities of LLMs requires premium quality prompts Eric (2022); Chen et al. (2023); Gajula (2023). Therefore, prompt engineering has attracted the attention of many researchers Varshney and Surla (2023); Meskó (2023); Wang (2023).
@@
Prompt engineering is a typically empirical science, which mainly involves designing and optimizing the prompts given to LLMs. As LLMs have the capability of natural language understanding, it is possible to ask them to execute tasks through unstructured natural language instructions directly. On this basis, some tricks for prompt optimization were first explored by researchers. Bsharat et al. (2023) introduce 26 guiding principles designed to make LLMs perform better. In addition to these directly usable prompt optimization tricks, some researchers have also focused on optimizing prompts based on historical data. Sun et al. (2023a) guide LLMs to derive new prompts for a given instance from the incorrect reasoning, and then summarise the corresponding prompts for each instance as a reference for optimizing the original prompts. Pryzant et al. (2023) leverage mini-batches of data to form natural language “gradients” and utilize beam search and bandit selection procedure to edit the current prompt in the opposite semantic direction of the gradient. Fan et al. (2023) analyze a large prompt database and present an automatic prompt optimization framework.
@@
Direct prompt optimization principles and methods based on historical data require a wealth of experience. Thus, these usually perform well only for specific tasks or domains. To improve generalization, some researchers have proposed adaptive prompt optimization methods. Guo et al. (2023) connect LLMs with evolutionary algorithms and proposes a novel framework for discrete prompt optimization, called EvoPrompt. Li et al. (2023a) design a multi-round dialogue alignment strategy and utilizes GPT-4 Achiam et al. (2023) to generation a readability prompt set. Meanwhile, they propose an efficient prompt screening metric that can filtrate high-quality prompts with linear complexity. Wang et al. (2023b) introduce PromptAgent which can reflect on model errors and generate constructive error feedback to induce precise expert-level insights and in-depth instructions. Hao et al. (2022) and Cheng et al. (2023) optimize prompts from the perspective of aligning human and LLMs’ preference styles.
@@
Prompt optimization can significantly improve the performance of LLMs, but the reusability of quality prompts is poor due to the lack of systematic design. Consequently, some researchers have devised rules for the construction of prompts. Nigh (2023) collects a large number of quality prompts and summaries the CRISPE rule for prompt design. Zamfirescu-Pereira et al. (2023) take a prototype LLM-based chatbot design tool as the design probe, supporting non-AI-experts engage in “end-user prompt engineering”. Some researchers have designed prompt construction rules for applications in different domains. Cao et al. (2023) present various prompt templates on deep learning program repair tasks for ChatGPT. Yeh et al. (2022) reformulate the biomedical relation extraction task as a cloze-test task under a simple prompt formulation to systematically generate comprehensive prompts. Liu and Chilton (2022) have evaluated 5493 generations covering 51 themes and 51 styles throughout five experiments in the text-to-image task and summaries prompt design guidelines.
@@
These methods, which are based on a great deal of experience in use, are mainly based on listed design rules and lack systematicity. In addition, these rules have strong domain relevance and model relevance, with low generalisability, flexibility, and reusability. To further unleash the performance of LLMs, some researchers have defined agents. Agents empower LLMs to use tools, acquire domain knowledge, retain long-term or short-term memories, and plan Xu et al. (2023); Xi et al. (2023); Park et al. (2023). Although agent methods Chase (2022); Hong et al. (2023); Wu et al. (2023) have systematically designed the key components of the prompt and reserved flexible customization interfaces, the learning costs are very high. In addition, it is difficult for non-AI experts to modify agent designs, so excellent designs are less reusable.
@@
To promote LLM-based applications and further stimulate the potential of LLMs, we would like to design a high-quality, reusable prompt template. This template should be extensible and generalizable. In addition, the template ought to be easy to learn and easy to work with. Inspired by the belief that prompt is the programming language of the LLM era Alouani (2023); Mund (2023), we have designed Language for GPT-like LLMs (LangGPT), a prompt design framework as the programming language for LLMs. LangGPT refers to the systematic, prescriptive, and reusable properties of programming languages, and retains the flexibility and extensibility of natural languages. We analyzed the differences between natural languages and programming languages to determine the properties that the prompt template should have. LangGPT is designed as a dual-layer structure, consisting of modules and internal elements. Modules in LangGPT can be divided into two categories: inherent modules and extension modules. For inherent modules, we design the necessary internal elements of each module in detail and give example templates. In addition, for the extension modules, we unified the design of the basic internal elements. Experiments have demonstrated that LangGPT is better than baseline prompts for bootstrapping LLMs. Furthermore, LLMs can leverage this framework to automatically generate prompts as if generating codes.
In summary, the contributions of this work include:
  • We proposed a dual-layer structured prompt design framework LangGPT to improve the generalization and reusability of prompts. In addition, we wrote detailed rules for designing prompts based on LangGPT, reducing the learning cost of prompt design and enabling LLMs to generate high-quality prompts automatically.
  • We proved experimentally that the prompts designed based on LangGPT can better guide LLMs to execute tasks. Meanwhile, we exemplified that LangGPT can help LLMs generate high-quality prompts.
  • We have built an online community based on LangGPT, which provides usage documentation and prompt design interfaces. In addition, collecting and sharing excellent prompt cases promotes the exchange of LLM applications. We conducted a user experience survey in the community to verify the ease of use and reusability of LangGPT.
2 Prompt Design Rules with Programming Language
@@
Programming languages are more standardized and reusable compared to natural languages. To design high-quality reusable prompts, we analyzed the differences between natural languages and programming languages and proposed prompt design rules.
@@ 2.1 Difference between Programming Languages and Natural Languages
While natural languages are primarily used for communication, programming languages are designed to define instructions for machines to execute tasks GeeksforGeeks (2023). The different purposes of application have led to very different contexts for the creation and evolution of the two languages.
  • 1.
    Target audience. Natural languages are spoken by humans to humans Grune et al. (2012), while programming languages are prepared by humans for machines Chakray (2018).
  • 2.
    Structure. The fact that computers can only understand fixed instructions requires programming languages to have strict, rigorous syntax and semantics. In opposition, natural languages have loose and flexible syntax and semantics, permitting creativity and variation and possessing a high degree of fault tolerance.
  • 3.
    Ambiguity. Natural languages are more ambiguous, but humans have the ability to clarify the meaning of expressions. For example, problems such as jumbled order and typos have less of an impact on reading, and mispronunciations can be understood Chakray (2018); Aho (2007). Programming languages are less ambiguous because they need to provide computers with clear instructions GeeksforGeeks (2023).
  • 4.
    Evolution and updating. Natural languages evolve naturally over time through human use and communication Sipser (1996). Natural languages are flexible in adding new words and meanings and discarding outdated usages Fromkin et al. (2018). Programming languages are specifically designed to communicate with machines Sebesta (2012). New syntax rules and features require explicit upgrades or releases Pratt et al. (1984).
@@
2.2 Rules for Prompt Design
We improve prompts by referring to the design ideas of programming languages. After analyzing the differences between programming languages and natural languages, we propose the design rules for prompts: (1) Prompts should have a regularised format. Flexible and ambiguous natural languages are difficult to understand for LLMs. Format-constrained prompts make users’ purpose and requirements more salient. (2) The structure of prompts should be extensible. Custom structures facilitate users to design suitable prompts according to their domains and tasks. (3) Specific requirements must be clear and complete. Both instructions and additional requirements should be explicit and complete to avoid misunderstandings or biases. (4) Languages should be flexible. Where the requirements are clear, flexible languages can be better adapted to different domains. In addition, flexible languages are easy for to learn users.
@@ 3 Natural Language Programming Framework of Prompts for LLMs
Based on the design rules, we present LangGPT, the natural language programming framework with a dual-layer architecture for LLMs.
@@ 3.1 Overall Dual-layer Structure
To systematically design prompts that meet the rules, we have made full reference to the design ideas and structures of object-oriented programming languages Rentsch (1982); Lutz (2010). We consider the prompt as a software project and analogize the prompt design process with the software development process. The correspondence is shown in Figure 1.
Based on analogical analyses, it can be found that natural language prompts have a similar multi-level structure as programming languages. Therefore, we refer to the structure of programming languages and propose a dual-layer structure for prompt design, and define the notion of module and element for prompts.
@@
A dual-layer structure can be a good way to standardize the formatting of prompts. However, the flexibility of natural languages would be lost if prompts are too strictly required to follow predefined standard modules and internal elements. In addition, it will reduce the generalisability of LangGPT to different tasks in different domains, which is not conducive to the reuse of quality prompts. To solve these problems, we divided the types of modules and elements into prompts. We defined inherent module and basic element as the predefined dual-layer prompt template. In addition, we constructed extension module and custom element that support customization. We provide both Markdown Gruber (2012) and JSON Pezoa et al. (2016) formats for inherent modules and extension modules. Furthermore, we have written basic elements for different modules and defined rules for writing custom elements.
@@ 3.3 Internal Basic Elements
[…]
Module Samples of Basic Elements
Profile
$\bullet$ You are a magazine editor.
Goal
$\bullet$ You need to generate a title for the article.
Constraint
$\bullet$ The length of the title should not exceed 20 words.
Workflow
### Extracting the kernel content
[…]
$\circ$ Analyse the theme of the article;
$\circ$ Summarising the core content from the article;
$\circ$ Save the kernel content.
Style
$\bullet$ The style of the title should be formal.
Table 2: Examples of basic internal elements of inherent modules in the writing scenario. This prompt belongs to the writing category and its purpose is to generate a title for a given article. We have chosen five modules as examples - profile, goal, constraint, workflow, and style - and show one internal element from each module. Particularly, for the workflow module, we show a function-like basic element.
Figure 2: Design process for extension modules and custom elements. For application scenarios where inherent modules cannot cover all aspects, new extension modules can be defined as needed. After defining the extension module, the internal elements should also be designed according to the requirements of the extension module. In addition, if inherent modules are compatible with the needs of the scenario but the basic elements cannot meet all of the requirements, custom elements can be added directly to the inherent modules. It is important to note that extension and inherent modules should be as mutually exclusive as possible to minimize modification costs.
@@ 3.4 Extension Module and Custom Element
Therefore, in addition to inherent modules and basic elements, we define the extension module and custom element to improve the generalization and reusability of prompts. The design flow of extension modules and custom elements is shown in Figure 2.
This process helps users quickly analyze their needs and come up with the key points that need to be included in the prompt. Based on the design templates provided by LangGPT and this design process, quality prompts can be efficiently designed to match the application scenarios.

### v2

@@ Abstract
LLMs have demonstrated commendable performance across diverse domains. Nevertheless, formulating high-quality prompts to instruct LLMs proficiently poses a challenge for non-AI experts. Existing research in prompt engineering suggests somewhat scattered optimization principles and designs empirically dependent prompt optimizers. Unfortunately, these endeavors lack a structured design template, incurring high learning costs and resulting in low reusability. In addition, it is not conducive to the iterative updating of prompts. Inspired by structured reusable programming languages, we propose LangGPT, a dual-layer prompt design framework as the programming language for LLMs. LangGPT has an easy-to-learn normative structure and provides an extended structure for migration and reuse. Experiments illustrate that LangGPT significantly enhances the performance of LLMs. Moreover, the case study shows that LangGPT leads LLMs to generate higher-quality responses. Furthermore, we analyzed the ease of use and reusability of LangGPT through a user survey in our online community [^1] The homepage: https://langgpt.ai, the Github project: https://github.com/langgptai/LangGPT, the tools and experiments: https://github.com/sci-m-wang/LangGPT-tools, and the Huggingface organization: https://huggingface.co/langgptai.
LangGPT: Rethinking Structured Reusable Prompt Design Framework for LLMs from the Programming Language
Ming Wang1, Yuanzhong Liu2, Xiaoyu Liang3, Songlian Li2, Yijie Huang1, Xiaoming Zhang1, Sijia Shen4, Chaofeng Guan4, Daling Wang1, Shi Feng1, Huaiwen Zhang1, Yifei Zhang1, Minghui Zheng3, Chi Zhang1, 1Northeastern University, 2Wuhan University, 3Zhejiang University, 4Shenyang Aerospace University Correspondence: wangdaling@cse.neu.edu.cn, liu.yuanzhong@foxmail.com
@@ 1 Introduction
Large language models (LLMs) can execute diverse tasks (Sun et al., 2023b, c; Yu et al., 2023) based on powerful language comprehension, reasoning, and generation capabilities. Injecting domain knowledge also enables LLMs to perform domain-related specific tasks (Wang et al., 2023a; Li et al., 2023c). Thus, LLM-based applications have attracted the attention of many non-AI experts, who have shown keen interest in using LLMs for programming, writing articles, medical diagnostics, etc (Zamfirescu-Pereira et al., 2023a). However, fully unleashing these capabilities of LLMs requires premium quality prompts (Eric, 2022; Chen et al., 2023; Gajula, 2023), but non-AI experts, even some AI-related practitioners, may not have the experience to write such prompts. Therefore, prompt engineering has been widely focused on (Varshney and Surla, 2023; Meskó, 2023; Wang, 2023).
@@
Prompt engineering is a typically empirical science, which mainly involves designing and optimizing the prompts. As LLMs can understand natural language, it is possible to ask them to execute tasks through unstructured natural language instructions directly. On this basis, some tricks for prompt optimization were first explored by researchers. Bsharat et al. (2023) introduce 26 guiding principles designed to improve the performance of LLMs. Liu and Chilton (2022) have evaluated 5493 generations covering 51 themes and 51 styles throughout five experiments in the text-to-image task and summaries prompt design guidelines. In addition to these tricks, some researchers have also focused on optimizing prompts based on historical data. Sun et al. (2023a) guide LLMs to derive new prompts for a given instance from the incorrect reasoning, and then summarise the corresponding prompts for each instance as a reference for optimizing prompts. Pryzant et al. (2023) leverage mini-batches of data to form natural language “gradients” and utilize beam search and bandit selection procedure to edit the current prompt in the opposite semantic direction of the gradient. Fan et al. (2023) analyze a large prompt database and present an automatic prompt optimization framework.
@@
Direct prompt optimization principles and methods based on historical data require a wealth of experience. Moreover, these usually perform well only for specific tasks or domains. To improve generalization, some researchers have proposed adaptive prompt optimization methods. Guo et al. (2023) connect LLMs with evolutionary algorithms and proposes a novel framework for discrete prompt optimization, called EvoPrompt. Li et al. (2023a) design a multi-round dialogue alignment strategy and utilizes GPT-4 (Achiam et al., 2023) to generate a readability prompt set. Meanwhile, they propose an efficient prompt screening metric that can filtrate high-quality prompts with linear complexity. Wang et al. (2023b) introduce PromptAgent which can reflect on model errors and generate constructive error feedback to induce precise expert-level insights and in-depth instructions. Hao et al. (2022) and Cheng et al. (2023) optimize prompts by aligning human and LLMs’ preference styles.
@@
Prompt optimization can significantly improve the performance of LLMs. However, such fragmented, unstructured prompt optimization principles lead to low reusability of quality prompts. In addition, it is not conducive to iterative updating of prompts. Consequently, some researchers have devised rules for the construction of prompts. Nigh (2023) collects vast quality prompts and summaries of the CRISPE rule for prompt design. Zamfirescu-Pereira et al. (2023b) take a prototype LLM-based chatbot design tool as the design probe, supporting non-AI-experts engage in “end-user prompt engineering”. Some researchers have designed prompt construction rules for applications in different domains. Cao et al. (2023) present various prompt templates on deep learning program repair tasks for ChatGPT. Yeh et al. (2022) reformulate the biomedical relation extraction task as a cloze-test task under a simple prompt formulation.
@@
These methods, which are based on a great deal of experience in use, have strong domain relevance and model relevance, with low generalisability, flexibility, and reusability. To further unleash the performance of LLMs, some researchers have defined agents. Agents empower LLMs to use tools, acquire domain knowledge, retain long-term or short-term memories, and plan (Xu et al., 2023; Xi et al., 2023; Park et al., 2023). Although agent methods (Hong et al., 2023; Wu et al., 2023) have systematically designed the key components of the prompt and reserved flexible customization interfaces, the learning costs are very high. In addition, it is difficult for non-AI experts to modify agent designs.
Aspects Natural Language Programming Language
Audience Natural languages are spoken by humans to humans (Grune et al., 2012). Programming languages are prepared by humans for machines (Chakray, 2018).
Structure Natural languages have loose and flexible syntax and semantics, permitting creativity and variation and possessing a high degree of fault tolerance. Computers can only understand fixed instructions and require programming languages to have strict, rigorous syntax and semantics.
Ambiguity Natural languages are more ambiguous, but humans can clarify the meaning of expressions (Chakray, 2018; Aho, 2007). Programming languages are less ambiguous because they need to provide computers with clear instructions (GeeksforGeeks, 2023).
Evolution Natural languages evolve naturally over time through human use and communication (Sipser, 1996). Natural languages are flexible in adding new words and meanings and discarding outdated usages (Fromkin et al., 2018). Programming languages are specifically designed to communicate with machines (Sebesta, 2012). New syntax rules and features require explicit upgrades or releases (Pratt et al., 1984).
Table 1: Difference between programming languages and natural languages
@@
To promote LLM-based applications and further stimulate the potential of LLMs, we would like to design a generalizable, reusable, and extensible prompt template. In addition, the template ought to be easy to learn and easy to work with for non-AI experts. Inspired by the belief that prompt is the programming language of the LLM era (Alouani, 2023; Mund, 2023), we have designed Language for GPT-like LLMs (LangGPT), a prompt design framework. LangGPT refers to the systematic, prescriptive, and reusable properties of programming languages, and retains the flexibility and extensibility of natural languages. We analyzed the differences between natural languages and programming languages and designed LangGPT as a dual-layer structure, which consists of modules and internal elements. Modules in LangGPT can be divided into two categories: inherent modules and extension modules. For inherent modules, we design the necessary internal elements of each module in detail and give example templates. For the extension modules, we unified the design of the basic internal elements. Experiments and case studies have demonstrated that LangGPT is better than baseline prompts for bootstrapping LLMs. Furthermore, with the belief that better instruction understanding leads to greater performance of LLMs (Ouyang et al., 2022; Chung et al., 2024), we constructed a dataset for LangGPT prompt generation. Experiments demonstrate that LLMs trained on the LangGPT prompt tasks have significantly improved performance. In addition, we conducted user research to prove the ease of use and user satisfaction of LangGPT based on our community.
In summary, the contributions of this work include: (1) We proposed a dual-layer structured prompt design framework LangGPT to improve the generalization and reusability of prompts, reducing the learning cost of prompt design. (2) We demonstrated through experiments and case studies that LangGPT prompts can better guide LLMs in executing tasks. Meanwhile, we experimentally proved that training on generating LangGPT prompts can improve the performance of LLMs. (3) We have built an online community based on LangGPT and conducted a user survey to verify the ease of use and reusability of LangGPT.
2 Prompt Design Principles with Programming Language
@@
Programming languages are more standardized and reusable compared to natural languages. To design high-quality reusable prompts, we analyzed the differences between natural languages and programming languages and proposed prompt design principles.
@@ 2.1 Difference between Programming Languages and Natural Languages
While natural languages are primarily used for communication, programming languages are designed to instruct machines to execute tasks (GeeksforGeeks, 2023). The different purposes have led to very different contexts for the creation and evolution of the two languages. The main differences between these languages are shown in Table 1.
@@
2.2 Principles for Prompt Design
We improve prompts by referring to the design ideas of programming languages. After analyzing the differences between programming languages and natural languages, we propose the design principles for prompts: (1) Prompts should have a regularised format. Flexible and ambiguous natural languages are difficult to understand for LLMs. Format-constrained prompts make users’ purpose and requirements more salient. (2) The structure of prompts should be extensible. Custom structures facilitate users to design suitable prompts according to their domains and tasks. (3) Specific requirements must be clear and complete. Both instructions and additional requirements should be explicit and complete to avoid misunderstandings or biases. (4) Languages should be flexible. Where the requirements are clear, flexible languages can be better adapted to different domains. In addition, flexible languages are easy for users to learn.
@@ 3 Natural Language Programming Framework of Prompts for LLMs
Based on the design principles, we present LangGPT, the natural language programming framework with a dual-layer architecture for LLMs.
@@ 3.1 Overall Dual-layer Structure
To systematically design prompts that meet the principles, we have made full reference to the design ideas and structures of object-oriented programming languages Rentsch (1982); Lutz (2010). We consider the prompt as a software project and analogize the prompt design process with the software development process. The correspondence is shown in Figure 1.
Based on analogical analyses, it can be found that natural language prompts have a similar multi-level structure as programming languages. Therefore, we refer to the structure of programming languages propose a dual-layer structure for prompt design, and define the notion of module and element for prompts.
@@
A dual-layer structure can be a good way to standardize the formatting of prompts. However, the flexibility of natural languages would be lost if prompts are too strictly required to follow predefined standard modules and internal elements. In addition, it will reduce the generalisability of LangGPT to different tasks in different domains, which is not conducive to the reuse of quality prompts. To solve these problems, we divided the types of modules and elements into prompts. We defined inherent module and basic element as the predefined dual-layer prompt template. Moreover, we constructed extension module and custom element that support customization. We provide both Markdown Gruber (2012) and JSON Pezoa et al. (2016) formats for inherent modules and extension modules. Furthermore, we have written basic elements for different modules and defined principles for writing custom elements.
Module Samples of Basic Elements
Profile $\bullet$ You are a magazine editor.
Goal $\bullet$ You need to generate a title for the article.
Constraint $\bullet$ The length of the title should not exceed 20 words.
Workflow ### Extracting the kernel content
$\bullet$ For the given article articledelimited-$\langle\textsc{article}\rangle$⟨ article ⟩, please execute the following actions:
   $\circ$ Analyse the theme of the article;
   $\circ$ Summarising the core content from the article;
   $\circ$ Save the kernel content.
Style $\bullet$ The style of the title should be formal.
Table 3: Examples of basic internal elements of inherent modules in the writing scenario. This prompt leads LLMs to generate a title for a given article. We have chosen five modules as examples i.e. profile, goal, constraint, workflow, and style. Particularly, for the workflow module, we show a function-like element.
@@ 3.3 Internal Basic Elements
The first of these is very similar to the assignment of properties or variables in programming languages. Correspondingly, the last two categories are similar to functions in programming languages. Thus, we construct these three types of basic elements. We use “The propertydelimited-$\langle\textsc{property}\rangle$⟨ property ⟩ is valuedelimited-$\langle\textsc{value}\rangle$⟨ value ⟩.” to simulate an assignment. For the latter two cases, it is necessary to specify the input information, the task, and the output, where input and output can be omitted. We simulate functions using a form like this: “For the given propertydelimited-$\langle\textsc{property}\rangle$⟨ property ⟩ of valuedelimited-$\langle\textsc{value}\rangle$⟨ value ⟩, please execute the following actions: actionsdelimited-$\langle\textsc{actions}\rangle$⟨ actions ⟩; Return the resultdelimited-$\langle\textsc{result}\rangle$⟨ result ⟩.” In the basic element writing patterns we provide, the contents contained in the angle brackets need to be populated according to the module and the usage scenario. It is important to note that the writing patterns we have provided only specify the idea of writing internal elements. To improve the generalisability and flexibility of prompts, the language can be adapted to express key information. In Table 3 we show some examples of the basic elements in some modules.
@@ 3.4 Extension Module and Custom Element
Therefore, in addition to inherent modules and basic elements, we define templates of extension modules and custom elements to improve the generalization and reusability of prompts. For application scenarios where inherent modules cannot cover all aspects, new extension modules can be defined as needed. After defining the extension module, the internal elements should also be designed according to the requirements of the extension module. In addition, if inherent modules are compatible with the needs of the scenario but the basic elements cannot meet all of the requirements, custom elements can be added directly to the inherent modules.